OpenStudy (anonymous):

Find the following. Assume that x can represent any real number. Use absolute value notation when necessary. sqrt9x^2

jimthompson5910 (jim_thompson5910):

\[\sqrt{9x^2}=\sqrt{9}\sqrt{x^2}=3|x|\]